OpenStudy (anonymous):

How many three-letter permutations can be made from the letters in the world ORANGE?

OpenStudy (amistre64):

5*4*3 = 20*3 = 60 maybe?

OpenStudy (amistre64):

ack!!... 6 letters to choose from lol

OpenStudy (amistre64):

6.5.4 = 120

OpenStudy (amistre64):

am i right?

OpenStudy (anonymous):

would you multiply 6x5x4x3x2x which woud =720 then multiply 3x2x1 andget 6 and divide 6 into 720? so the final answe woul b 420.